# Infrastructure for AI Agents

### v1

@@ 4.3 Inter-Agent Communication
Existing alternatives: Existing communication infrastructure may be difficult for agents to use because they require human involvement. For example, signing up for a Gmail account requires access to a phone number. It may also be difficult to add agent-specific functionality onto existing communication infrastructure. For instance, Facebook Messenger does not currently support broadcasting functions.
Adoption: The difficulties of using the existing alternatives above could spur adoption of more specialized inter-agent communication infrastructure.
Limitations: As with existing internet communication infrastructure, inter-agent communication infrastructure could be abused. For example, broadcast functionalities could be frequent sources of spam. Scammers Could use an addressing system to target agents they believe to be especially vulnerable (e.g., more easily hijacked through a prompt injection). The extent of abuse will likely depend upon the specifics of implementation (e.g., the effectiveness of any blocking functionalities).
@@
  • What security measures should inter-agent communication infrastructure incorporate? For example, what could be analogues to HTTPS?
  • What are use cases for broadcasting?
  • How should broadcasting be designed? For example, one way of implementing broadcasting is to rely on other agents to pass messages, rather than having a centralized system.
  • What functionalities should inter-agent communication infrastructure include? For example, a communication protocol could include a way to signal or carry out commitments (Section 4.4).
@@ 4.4 Commitment Devices
Existing alternatives: Commitment devices that depend upon (threats of) action in the physical world or human social norms, such as legal contracts or making announcements to friends, are likely to be ineffective for agents that we are currently building. Connecting agents to existing software-based commitment devices, such as smart contracts, is likely to be more fruitful in the short term.
Adoption: Commitment devices are low-cost for actors who seek to make conditional commitments. If the conditions of the commitment do not hold (e.g., a quorum is reached), the actors take no action. Some types of commitment devices, such as dominant assurance contracts (Tabarrok, 1998), may even positively incentivize participation in the device.
@@
  • What commitment devices could help agents from different users better cooperate on productive projects (e.g., software development)?
@@
  • If future agents are responsive to legal incentives, [^10] One possibility is to directly train agents that follow the law. would new types of commitment devices still be necessary?
@@ 5.1 Incident Reporting
Others have also proposed incident reporting systems, but without the focus on agents as a source of information (Costanza-Chock et al., 2022; Wei & Heim, 2024). More broadly, analogous systems include financial whistleblower programs for financial institutions (Securities and Exchange Commission, 2024), public reporting systems for vehicle offenses (Government of the United Kingdom, 2024), and cybersecurity incident reporting (Government of the United Kingdom, 2022). Another related idea is using weaker AI systems to supervise stronger AI systems (Greenblatt et al., 2024).
@@ 5.2 Rollbacks
  • How could a rollback mechanism for agents connect to existing methods for undoing actions (e.g., a credit card company undoing financial transactions)?
  • Could the availability of a rollback option to the user be made conditional upon some measure of the agent’s reliability?
@@
6 Common Challenges
We discuss challenges common not only to all types of agent infrastructure, but also infrastructure (e.g., for the Internet) more broadly.
@@ 6.1 Lack of Interoperability
Actors could adopt mutually incompatible forms of agent infrastructure. For example, different groups of developers and providers of digital services could implement ID systems that do not recognize IDs from competitor systems. An agent with a valid certification under one ID system may not be able to interact with entities that recognize another ID system. Although this lack of interoperability would make IDs less useful, it could benefit dominant players by shutting out potential competitors and creating a walled garden around users. Similar problems beset digital platforms (e.g., Facebook Messenger, WhatsApp, iMessage). Legislation such as the EU’s Digital Markets Act has encouraged more interoperability in these cases, but it remains to be seen whether and how similar rules could apply to the implementation of agent infrastructure.
6.2 Difficulty of Use
Users will likely not opt for infrastructure that is unintuitive to use or requires substantial setup. For example, users could simply ignore oversight layers that flag too many actions.
@@ 6.3 Lock-In
Widespread adoption of harmful or ineffective infrastructure could be difficult to reverse. Consider the network effects of interaction protocols. Moving away from a common protocol could mean losing important traffic. As an additional barrier, the benefits of moving away from the protocol could be realized only once a sufficient number of entities do so. Adoption of the Border Gateway Protocol (BGP) (Rekhter et al., 2006) ran into similar problems. Routing information on BGP is not verifiable: any entity can claim to have a route to another entity. In 2008, in an attempt to block YouTube domestically, a Pakistani ISP accidentally sent routing information that brought down YouTube worldwide (Mathew, 2014). Despite the potential for such problems, adoption of a verifiable version of BGP has been slow (Testart et al., 2024). For agents, it will be important to adopt procedures for effectively updating infrastructure.
@@ 7 Related Work
Some recent literature has grown around how to govern foundation-model-based agents. Chan et al. (2023); Lazar (2024); Gabriel et al. (2024); Kolt (2024) study the ethical and societal implications of agents. Shavit et al. (2023) propose practices for governing agents, some of which are related to types of agent infrastructure we discuss. Chan et al. (2024a; b) propose agent infrastructure that can help certain actors obtain visibility into how agents are used, whereas we take a broader view of the functions of agent infrastructure. Finally, Perrier (2025); Perrier & Lazar (2025) study the ontology of agents and its relation to agent infrastructure.

### v2

@@ 4.3 Inter-Agent Communication
Existing alternatives: Some existing communication infrastructure may be difficult for agents to use because they require human involvement. For example, signing up for a Gmail account requires access to a phone number. It may also be difficult to add agent-specific functionality onto existing communication infrastructure. For instance, Facebook Messenger does not currently support broadcasting functions.
On the other hand, existing open protocols could be the foundation of more specific agent protocols. For example, Marro et al. (2024) describe a meta protocol wherein LLM agents can negotiate, in natural language, the subsequent use of more efficient, structured data formats (e.g. JSON). As another example, Google’s A2A protocol uses HTTP for transport between client and server agents (Surapaneni et al., 2025).
Adoption: At a high level, deficiencies in existing communication protocols could spur the adoption of more specialized alternatives for agents. As discussed above, required human intervention for existing communication infrastructure could slow or inconvenience interactions. On the other hand, since many agents are able to communicate in natural language, newer protocols could take advantage of this flexibility (Marro et al., 2024).
As we discuss further in Section 6.1, network effects are likely to be important for the adoption of specific communication protocols as common standards. Buy-in from agent developers and deployers, digital platforms, and large organizations that use agents is likely to be crucial. For example, Google is collaborating with the number of large enterprises for its A2A protocol (Surapaneni et al., 2025).
Limitations: As with any communication infrastructure, inter-agent communication infrastructure could be abused. For example, broadcast functionalities could be sources of spam. Scammers could use an addressing system (e.g., a “phone number” or “URL” for each agent) to target particular agents. Worms that target the adversarial vulnerabilities of AI agents (Cohen et al., 2024) could spread along communication channels. The severity of such problems will likely depend the development of countermeasures, such as whether particular agents can be blocked. Analogously, email scanners can help block malicious payloads and social engineering attempts.
More broadly, organizations could build communication protocols (and infrastructure in general) in ways that are beneficial to their financial interests but are detrimental to other parties. For example, ad-supported businesses could prefer communication protocols that make it difficult for agents to block advertising from other agents. While advertising may not be undesirable in and of itself, the dominance of ad-supported businesses could make it difficult for alternative business models to arise (Lazar et al., 2024).
@@
  • Beyond encryption, how could inter-agent communication protocols be secured? For example, is there a way of making sure that messages do not contain jailbreaks or steganographic communications?
  • What are use cases for broadcasting, and how should it be designed?
  • What other functionalities should inter-agent communication infrastructure include? For example, a communication protocol could include a way to signal or carry out commitments (Section 4.4).
@@ 4.4 Commitment Devices
Existing alternatives: Many existing commitment devices depend upon (threats of) action in the physical world or human social norms, such as legal contracts or making public announcements. These threats or norms can be communicated to language-model-based agents. However, it is not clear how reliably such agents will respond. As one piece of evidence, Li et al. (2023) find that emotional cues can affect language model performance. On the other hand, language models can still fail to reliably follow instructions (Baker et al., 2025).
Software-based commitment devices, such as smart contracts, are likely to be more fruitful in the short term. For example, agent $A$ could commit to transfer an amount under certain conditions to agent $B$ simply by writing a smart contract and deploying it onto the blockchain, which lets agent $B$ verify the contract. This use case is no different from if two humans want to commit to the same agreement. In this sense, agent-specific modifications to the core technology of smart contracts are not necessarily needed. Rather, the main limitation is the types of transactions that can be carried out on a blockchain. Many agreements involve an information source in the real world. For example, an agreement about a building might involve terms related to its appearance or other physical conditions. Without a reliable digital connection to such sources, smart contracts are inapplicable.
Adoption: Since commitment devices facilitate agreements, there is a market incentive to build and adopt commitment devices. Analogously, many private companies provide escrow services. Kickstarter is an example of an assurance contract.
Adoption of commitment devices that depend upon a blockchain could face barriers. One potential barrier is simply not many types of transactions cannot yet be carried out on a blockchain, as discussed above. Another barrier is potential lack of consumer trust in blockchain technologies, given high-profile scandals (Sherman & Hoskins, 2023).
@@
  • How can adoption of software-based commitment devices, such as smart contracts, be improved?
@@
  • If future agents are responsive to legal incentives, [^10] One possibility is to directly train agents that follow the law (O’Keefe et al., 2025) what other commitment devices would still be necessary?
@@ 5.1 Incident Reporting
Others have also proposed incident reporting systems, but without the focus on agents as a source of information (Costanza-Chock et al., 2022; Wei & Heim, 2024). More broadly, analogous systems include financial whistleblower programs for financial institutions (Securities and Exchange Commission, 2024), public reporting systems for vehicle offenses (Government of the United Kingdom, 2024), and cybersecurity incident (Government of the United Kingdom, 2022) or vulnerabilities (MITRE Corporation, 2025) reporting. Another related idea is using weaker AI systems to supervise stronger AI systems (Greenblatt et al., 2024).
@@ 5.2 Rollbacks
Another potential issue is moral hazard. If rollbacks were readily available, developers could have weaker incentives to implement effective oversight layers. Similarly, users could have weaker incentives to properly monitor their agents. The cost of oversight failures would be transferred to those operating such rollback mechanisms. A potential way to deal with moral hazard is through insurance terms. For example, rollbacks could be made available only if users pay on insurance premium, deductibles, and/or co-payments.
@@
  • Could access to a rollback mechanism be made conditional upon some measure of the agent’s reliability?
  • How could rollbacks interact with insurance?
@@
6 Challenges and Limitations
We discuss challenges and limitations common to many types of agent infrastructure.
6.1 Adoption
A common challenge for all infrastructure is adoption. Some infrastructure mostly depends upon agent developers or deployers adopting it, such as oversight layers. Other infrastructure depends upon coordination between multiple parties. For example, two parties can only communicate if they use the same communication protocol. In these cases, network effects are likely to be crucial. Infrastructure is more likely to be widely adopted if the creators can take advantage of existing networks or build them. For example, Google has partnered with a number of large enterprises in the introduction of its recent agent-agent communication protocol (Surapaneni et al., 2025). Smaller and less well-known players could be less able to compete, even if their infrastructure offers superior features (also see Section 6.2). In addition to actors that build agent infrastructure, existing internet standards organizations could also be crucial for diffusion. Indeed, much agent infrastructure will build upon existing internet protocols, and some AI agent working groups already exist. [^11] E.g., see https://www.w3.org/groups/cg/webagents/
@@ 6.2 Lack of Interoperability
Infrastructure from different developers could be mutually incompatible. For example, an ID system could fail to recognize IDs (or certifications) from competitor systems. Although this lack of interoperability would make IDs less useful, it could benefit dominant players by shutting out potential competitors. For example, an agent developer could try to promote use of its own agents by limiting the interoperability of its reputation system. Analogously, a restaurant cannot transfer its profile from Yelp to Google. Similar problems beset other digital platforms (e.g., Facebook Messenger, WhatsApp, iMessage). Legislation such as the EU’s Digital Markets Act has encouraged more interoperability in these cases, but it remains to be seen whether and how similar rules could apply to the implementation of agent infrastructure.
@@ 6.3 Lock-In
Widespread adoption of insufficiently secure or otherwise harmful infrastructure could be difficult to reverse. Consider the network effects of interaction protocols. Moving away from a common protocol could mean losing important traffic. As an additional barrier, the benefits of moving away from the protocol could be realized only once a sufficient number of entities do so. Adoption of the Border Gateway Protocol (BGP) (Rekhter et al., 2006) ran into similar problems. Routing information on BGP is not verifiable: any entity can claim to have a route to another entity. In 2008, in an attempt to block YouTube domestically, a Pakistani ISP accidentally sent routing information that brought down YouTube worldwide (Mathew, 2014). Despite the potential for such problems, adoption of a verifiable version of BGP has been slow (Testart et al., 2024). For agents, it will be important to adopt procedures for effectively updating infrastructure.
@@ 7 Related Work
Past work has considered the protocols needed to create and manage networks of agents (Kahn & Cerf, 1988; Poslad, 2007; Wooldridge, 2009). While the specific details of such protocols may not be applicable to how today’s agents operate, conceptual ideas such as a communication system between agents (Kahn & Cerf, 1988) are still relevant. Our paper builds upon this prior work by accounting for (1) unique features of more general-purpose, foundation-model-based agents and (2) modern digital infrastructure that could support agent infrastructure. Furthermore, we describe in detail how agent infrastructure can help to manage the risks of agents.
Some recent literature focuses on how to govern foundation-model-based agents. Chan et al. (2023); Lazar (2024); Gabriel et al. (2024); Kolt (2024) study the ethical and societal implications of agents. Shavit et al. (2023) propose practices for governing agents, some of which are related to types of agent infrastructure we discuss. Chan et al. (2024a; b) propose agent infrastructure that can help certain actors obtain visibility into how agents are used, whereas we take a broader view of the functions of agent infrastructure. Hammond et al. (2025) propose a taxonomy of risks from multi-agent systems. Finally, Perrier (2025); Perrier & Lazar (2025) study the ontology of agents and review historical examples of agent infrastructure.